OpenStudy (zeig_101):

Can someone teach me how to do this sort of problem? Let f(x)=-5x-4 and g(x)=6x-7. Find f(x)+g(x)

OpenStudy (anonymous):

stick a plus sign between them

OpenStudy (anonymous):

after that, combine like terms

OpenStudy (zeig_101):

So f(x)=-5x-4+g(x)=6x-7?

OpenStudy (anonymous):

\[-5x-4+6x-7\]

OpenStudy (zeig_101):

OH ok, so -5x-4+6x-7 Then X-11

OpenStudy (anonymous):

\[\overbrace{-5x-4}^{f(x)}+\overbrace{6x-4}^{g(x)}\]

OpenStudy (zeig_101):

And that's my answer?

OpenStudy (anonymous):

typo there, i meant \[\overbrace{-5x-4}^{f(x)}+\overbrace{6x-7}^{g(x)}\]

OpenStudy (zeig_101):

So it's x-11?

OpenStudy (anonymous):

but yes, you are correct \(f(x)+g(x)=x-11\)

OpenStudy (zeig_101):

Ok. There's also one more type of question, can I get help on it?

OpenStudy (anonymous):

sure

OpenStudy (zeig_101):

f(x)=x^2-16 and g(x)=4+4. find f/g and its domain. This one has me thoroughly confused.

OpenStudy (anonymous):

put a fraction bar between them

OpenStudy (anonymous):

i assume there is a typo in your question

OpenStudy (zeig_101):

So (x^2-16)/(x+4)?

OpenStudy (anonymous):

yes, that is the first step then some algebra

OpenStudy (anonymous):

before the algebra, make sure to find the domain since you cannot divide by zero, set \(x+4=0\) get \(x=-4\) and then say "all real numbers except \(-4\)"

OpenStudy (zeig_101):

x^2-16/(x+4=0) x=-4 -4^2-16/0 16-16/0 0/0? I think I did that wrong.

OpenStudy (anonymous):

yes, i think so you confused finding the domain with evaluating at \(x=-4\) you CANNOT

OpenStudy (anonymous):

evaluate at \(x=-4\) which is why the domain is all real numbers EXCEPT for \(-4\)

OpenStudy (zeig_101):

OH Ok, I think I get it

OpenStudy (anonymous):

you still have some algebra to do

OpenStudy (zeig_101):

x-4, and the domain is all real numbers except x=-4?

OpenStudy (anonymous):

yes

OpenStudy (zeig_101):

Okay. Thanks for the help!

OpenStudy (anonymous):

yw who does math on christmas eve? isn't it a holiday or something?

OpenStudy (zeig_101):

It is, but I like to work ahead in class so if I get sick or something I don't have to worry as much.

OpenStudy (anonymous):

good luck!